OpenStudy (surana):

Solve the inequality: (x+3) (x-5) < 0

OpenStudy (korosh23):

Are you looking for the value of the x?

OpenStudy (surana):

Possible answers being A: X < -3 or X > 5 B: X < -5 or X > 3 C: -3 < X < 5 D: -5 < X < 3

OpenStudy (korosh23):

Well it is easy! Lets go through it

OpenStudy (surana):

Alright, let's do that then.

OpenStudy (xapproachesinfinity):

draw a number line and locate x=- 3 and x=5 test some values before -3 and btw and after 5

OpenStudy (jdoe0001):

hint: \(\large { (x+3) (x-5) < 0 \implies \begin{cases} (x+3) < 0 \to x=?\\ (x-5) < 0 \to x=? \end{cases} }\)

OpenStudy (surana):

So I draw a line, find the numbers and see the values?

OpenStudy (xapproachesinfinity):

jdoe one positive one negative not all of them

OpenStudy (korosh23):

 tell me what is the value of ?